OpenStudy (lena772):

Arnold's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $40 a day for food and lodging and $0.50 for each mile traveled. Arnold drove 400 miles and was reimbursed $2600. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Arnold spend on this trip? (1 points)

ganeshie8 (ganeshie8):

food+lodging = $40 per day conveyance = $0.5 per mile

ganeshie8 (ganeshie8):

Since he drove 400 miles, how much he wud get for conveyance ?

OpenStudy (lena772):

240 total

OpenStudy (lena772):

200 for the conveyance

ganeshie8 (ganeshie8):

Yes, 200 for conveyance

ganeshie8 (ganeshie8):

say he was on trip for \(x\) days

ganeshie8 (ganeshie8):

for part A, can u setup an equation with the info above ?

ganeshie8 (ganeshie8):

if he was on trip for \(x\) days, then he wud get \(40x\) dollars

OpenStudy (lena772):

40x+200=2600

ganeshie8 (ganeshie8):

and since he got 200 $ for the conveyance, the equation wud be :- \(40x + 200 = 2600\)

OpenStudy (lena772):

:)

ganeshie8 (ganeshie8):

try part B, its tricky cuz u need to justify EACH step

OpenStudy (lena772):

40x+200=2600 Take 200 away from each side 40x+200-200=2600-200 You will then be left with 40x=2400 Divide by 40 on each side 40x/40=2400/40 You will then be left with x=60

ganeshie8 (ganeshie8):

looks good , but i dont see any algebraic properties used Part B: Solve this equation justifying each step with an \(\color{red}{algebraic ~~property}\) of equality. (6 points)

ganeshie8 (ganeshie8):

40x+200=2600 40x+200 - 200 = 2600 - 200 (By subtraction property of equality) 40x = 2400 40x/40 = 2400/40 (By division property of equality) x = 60

ganeshie8 (ganeshie8):

your teacher will be looking for those two property names :o

OpenStudy (lena772):

So subtraction can be first?

ganeshie8 (ganeshie8):

"subtraction property of equality"

OpenStudy (lena772):

Yes, but subtraction can go before division?

ganeshie8 (ganeshie8):

Yes why not

ganeshie8 (ganeshie8):

why do u ask that ?

ganeshie8 (ganeshie8):

40x+200=2600 when u have equation like above, subtracting 200 is the first thing u want to do. doing anything else is simply ridiculous

ganeshie8 (ganeshie8):

you can divide 40 both sides, but if u do that, u wil have to divide two times, one time for 200 and second time for 2600... so just do subtraction first

OpenStudy (lena772):

PEDMAS

ganeshie8 (ganeshie8):

we're still sticking to PEMDAS. but when solving equations, NEVER FOLLOW PEMDAS order.

ganeshie8 (ganeshie8):

40x+200=2600 you DONT need to divide 40 first

ganeshie8 (ganeshie8):

you can subtract 200 first, cuz its easy to do here

ganeshie8 (ganeshie8):

lets do it both ways, and see if we can convince ourself if we get both answers same ?

ganeshie8 (ganeshie8):

40x+200=2600 using PEMDAS order ``` 40x+200=2600 divide 40 both sides x + 5 = 65 subtract 5 both sides x = 60 ``` without using PEMDAS ``` 40x+200=2600 subtract 200 both sides 40x = 2400 divide 40 both sides x = 60 ```

OpenStudy (lena772):

Ok I'll do what we first did then :)

ganeshie8 (ganeshie8):

good :)